# Dynamical Lensing Tomography of Black Hole Ringdown

### v2

@@ A.2 Metric Reconstruction
    ^$\begin{split}\hat{\mathcal{E}}_{ab}(h)=\frac{1}{2}\left[-\nabla^{c}\nabla_{c}h_{ab}-\nabla_{a}\nabla_{b}h_{c}{}^{c}+2\nabla^{c}\nabla_{(a}h_{b)c}+g_{ab}\left(\nabla^{c}\nabla_{c}h_{d}{}^{d}-\nabla^{c}\nabla^{d}h_{cd}\right)\right]\,.\end{split}$ (17)
@@
  • The Teukolsky operator $\hat{\mathcal{O}}$, which can be written as Teukolsky (1973)
    $\hat{\mathcal{O}}_{0}=$ $\left(\hat{D}-3\epsilon+\bar{\epsilon}-4\rho-\bar{\rho}\right)\left(\hat{\Delta}-4\gamma+\mu\right)-\left(\hat{\delta}+\bar{\varpi}-\bar{\alpha}-3\beta-4\tau\right)\left(\hat{\bar{\delta}}+\varpi-4\alpha\right)-3\Psi_{2}\,,$ (18)
@@
    $\hat{\mathcal{O}}_{0}\Psi_{0}^{(1)}=8\pi T_{0}\,,\,\hat{\mathcal{O}}_{4}\Psi_{4}^{(1)}=8\pi T_{4}\,.$ (19)
    Their separable form, Teukolsky master equation, can be written as
    $\hat{\mathcal{O}}_{s}\psi_{s}=8\pi\Sigma T_{s}\,,$ (20)
@@
    $\psi_{-2}=\rho^{-4}\Psi_{4}^{(1)}\,,\hat{\mathcal{O}}_{-2}=2\Sigma\rho^{-4}\hat{\mathcal{O}}_{4}\rho^{4}\,,T_{-2}=2\rho^{-4}T_{4}\,;$ (21)
@@
    $\psi_{+2}=\Psi_{0}^{(1)},\,\hat{\mathcal{O}}_{+2}=2\Sigma\hat{\mathcal{O}}_{0}\,,T_{+2}=2T_{0}\,.$ (22)
  • The operator $\hat{\mathcal{S}}$, which acts on a energy-momentum tensor $T_{ab}$ to yield the source $T_{4}$ of the Teukolsky equation, (Eq. 2.15 in Ref. Teukolsky (1973))
    $\hat{\mathcal{S}}^{ab}=$ $\left(\hat{\Delta}+3\gamma-\bar{\gamma}+4\mu+\bar{\mu}\right)\left[(\hat{\overline{\delta}}-2\bar{\tau}+2\alpha)n^{(a}\bar{m}^{b)}-(\hat{\Delta}+2\gamma-2\bar{\gamma}+\bar{\mu})\bar{m}^{a}\bar{m}^{b}\right]$ (23)
@@
  • The operator $\hat{\mathcal{T}}$, which acts a metric perturbation $h_{ab}$ to the perturbed Weyl scalar $\Psi_{4}^{(1)}$.
    $\hat{\mathcal{T}}^{ab}=$ $-\frac{1}{2}\left\{\left(\hat{\overline{\delta}}-\bar{\tau}+3\alpha+\bar{\beta}\right)\left(\hat{\overline{\delta}}-\bar{\tau}+2\alpha+2\bar{\beta}\right)n^{a}n^{b}\right.$ (24)
@@
Figure 1: Relation of operators for $s=-2$.
As pointed out by Wald Wald (1978), one can find the following operator identity
$\hat{\mathcal{O}}_{4}\hat{\mathcal{T}}=\hat{\mathcal{S}}\hat{\mathcal{E}}\,,$ (25)
@@
^$\hat{\mathcal{E}}_{ab}(h)=2\operatorname{Re}(\hat{\mathcal{T}}^{\dagger}_{ab}\hat{\mathcal{O}}^{\dagger}_{4}\Phi)=0$ (26)
@@
$h_{ab}=2\operatorname{Re}\left(\hat{\mathcal{S}}_{ab}^{\dagger}\Phi\right)\,.$ (27)
Chrzanowski introduced two radiation gauges: the ingoing radiation gauge (IRG) and the outgoing radiation gauge (ORG) Chrzanowski (1975). In this paper, we choose ORG, which can be written as
$h_{ab}n^{b}=0=g^{ab}h_{ab}\,.$ (28)
@@
$\mathcal{O}^{\dagger}_{4}=\left(\hat{D}-3\epsilon+\bar{\epsilon}-\bar{\rho}\right)\left(\hat{\Delta}-4\gamma-3\mu\right)-\left(\hat{\delta}-3\beta-\bar{\alpha}+\bar{\varpi}\right)\left(\hat{\bar{\delta}}-4\alpha-3\varpi\right)-3\Psi_{2}\,.$ (30)
@@
$\hat{\mathcal{O}}_{4}^{\dagger}\Phi_{\mathrm{ORG}}=\rho^{-4}\mathcal{O}_{0}\left(\rho^{4}\Phi_{\mathrm{ORG}}\right)=0\,,$ (31)
and the metric component can be written as Dias et al. (2009)
$h_{ab}=$$\left\{n_{(a}\bar{m}_{b)}\left[\left(\hat{\bar{\delta}}+\bar{\beta}-3\alpha+\bar{\tau}+\varpi\right)\left(\hat{\Delta}-4\gamma-3\mu\right)\right.\right.$ (32)
@@
where “c.c.” stands for complex conjugate part of the whole object. If we define $\mathring{\Phi}_{\mathrm{ORG}}=\rho^{4}\Phi_{\mathrm{ORG}}$, then the non-zero metric component can be written as Lousto and Whiting (2002); van de Meent and Shah (2015)
$h_{ll}\equiv h_{\mu\nu}l^{\mu}l^{\nu}=\hat{\mathcal{H}}_{ll}^{\dagger}\mathring{\Phi}_{\mathrm{ORG}}+\text{c.c.}\,,\,h_{lm}\equiv h_{\mu\nu}l^{\mu}m^{\nu}=\hat{\mathcal{H}}_{lm}^{\dagger}\mathring{\Phi}_{\mathrm{ORG}},$ (33)
@@
^$\hat{\mathcal{H}}_{ll}^{\dagger}$ $=-\rho^{-4}\left(\hat{\bar{\delta}}-3\alpha-\bar{\beta}+5\varpi\right)\left(\hat{\bar{\delta}}-4\alpha+\varpi\right)\,,$ (34)
@@
While $\mathring{\Phi}_{\mathrm{ORG}}$ does indeed satisfy the vacuum Teukolsky equation with $s=+2$, the field does not correspond with either the Weyl scalar $\Psi^{\mathrm{(1)}}_{0}$ or $\Psi^{\mathrm{(1)}}_{4}$ originating from the reconstructed metric $h_{ab}$. From Fig. 1, it is clear that $\Psi^{\mathrm{(1)}}_{4}$ can be obtained via a fourth-order ordinary differential equation (ODE) along ingoing null rays Ori (2003); van de Meent and Shah (2015); Spiers et al. (2023)
$\Psi_{4}^{(1)}=\hat{\mathcal{T}}^{ab}h_{ab}\,.$ (35)
We can reduce this circularity condition to what is known as a radial inversion relation Spiers et al. (2023); Kofroň (2020); Aksteiner and Bäckdahl (2019); Deadman and Stewart (2010); Pound and Wardell (2020)
$\Psi_{4}^{(1)}=\frac{1}{2}\left(\hat{\Delta}+3\gamma-\bar{\gamma}\right)\left(\hat{\Delta}+2\gamma-2\bar{\gamma}\right)\left(\hat{\Delta}+\gamma-3\bar{\gamma}\right)\left(\hat{\Delta}-4\bar{\gamma}\right)\bar{\Phi}_{\mathrm{ORG}}\,,$ (36)
which can be solved in vacuum algebraically by applying the separation of variable in next section.
A.2.1 QNM to Hertz potential
In Boyer-Lindquist (BL) coordinates, the Teukolsky operator $\hat{\mathcal{O}}_{s}$ with spin $s$ in Eq. (20) can be written as Teukolsky (1973)
@@ A.2.1 QNM to Hertz potential
$\displaystyle-2s\left[\frac{M\left(r^{2}-a^{2}\right)}{\Delta_{\mathrm{BL}}}-r-ia\cos\theta\right]\frac{\partial\psi_{s}}{\partial t}+\left(s^{2}\cot^{2}\theta-s\right)\psi_{s}$ (37)
@@
$\psi_{s}(t,r,\theta,\phi)=\frac{1}{\sqrt{2\pi}}\sum_{l,m,n}{}_{s}Z_{\ell mn}\,{}_{s}S_{\ell mn}(\theta)\,{}_{s}R_{\ell mn}(r)e^{-i\omega_{mn}t+im\phi}\,,$ (38)
@@
$\partial_{z}\left[\left(1-z^{2}\right)\partial_{z}\left[{}_{s}S_{\ell mn}(\theta)\right]\right]+\left[(cz)^{2}-2csz+s+A^{\ell mn}_{s}(c)-\frac{(m+sz)^{2}}{1-x^{2}}\right]{}_{s}S_{\ell mn}(\theta)=0$ (39)
where $c=a\omega$ and ${}_{s}S_{\ell mn}(\theta)e^{im\phi}$ is the spin-weighted spheroidal harmonics. To normalize the spin-weighted spheroidal harmonics, we apply the following normalization condition
@@
In the following, we will revisit the procedure for obtaining the Hertz potential corresponding to specific Weyl scalar $\Psi^{\mathrm{(1)}}_{0}$ or $\Psi^{\mathrm{(1)}}_{4}$, which is discussed in Ref. Ori (2003); van de Meent and Shah (2015). According to the previously mentioned condition, $\hat{\mathcal{O}}_{+2}\mathring{\Phi}_{\mathrm{ORG}}=0$, we can expand $\mathring{\Phi}_{\mathrm{ORG}}$ in modes
$\mathring{\Phi}_{\mathrm{ORG}}=\frac{1}{\sqrt{2\pi}}\sum_{\ell,m,n}C_{\ell mn}\,{}_{+2}R_{\ell mn}(r)\,{}_{+2}S_{\ell mn}(\theta)e^{im\phi-i\omega_{mn}t}\,,$ (43)
where $C_{\ell mn}$ are constant coefficients. In addition, the Hertz potential $\mathring{\Phi}_{\mathrm{ORG}}$ also satisfies Eq. (36), which can be simplified using the following relationship
$\left(\hat{\Delta}-m\bar{\gamma}+n\gamma\right)f=\Delta_{\mathrm{BL}}^{(n-m)/2}\bar{\rho}^{-m}\rho^{n}\hat{\Delta}\left[\Delta_{\mathrm{BL}}^{(m-n)/2}\bar{\rho}^{m}\rho^{-n}f\right]\,.$ (44)
In this equation, $f=f(t,r,\theta,\phi)$ can represent any function. Then Eq. (36) can be written as Lousto and Whiting (2002); van de Meent and Shah (2015); Ori (2003); Pound and Wardell (2020)
$\frac{1}{32}\Delta_{\mathrm{BL}}^{2}\left(\hat{\mathcal{D}}_{0}^{\dagger}\right)^{4}\left[\Delta_{\mathrm{BL}}^{2}\bar{\mathring{\Phi}}_{\mathrm{ORG}}\right]=\rho^{-4}\Psi_{4}^{\mathrm{(1)}}\,.$ (45)
@@
$\hat{\mathcal{D}}_{0}^{\dagger}=\partial_{r}-\frac{\left(r^{2}+a^{2}\right)\partial_{t}+a\partial_{\phi}}{\Delta_{\mathrm{BL}}}\,.$ (46)
Note that Eq. (45) can be solved easily on the mode basis given by Eq. (43), using the operator
$\hat{\mathcal{D}}_{mn}^{\dagger}=\partial_{r}+i\frac{K_{mn}}{\Delta_{\mathrm{BL}}}=\partial_{r}+i\frac{\omega_{mn}\left(r^{2}+a^{2}\right)-ma}{\Delta_{\mathrm{BL}}}\,,$ (47)
which act on radial modes. The coefficient can be fixed by considering the asymptotic limit approaching infinity and the horizon. This analysis was previously carried out in Ref. Ori (2003); van de Meent and Shah (2015) for real $\omega$. Due to our choice of gauge (ORG), the results are similar to those in Section III.C of Ref. van de Meent and Shah (2015). Here, we briefly review and extend the procedure to complex $\omega$.
$\frac{1}{32}\Delta_{\mathrm{BL}}^{2}\left(\hat{\mathcal{D}}_{0}^{\dagger}\right)^{4}\left[\Delta_{\mathrm{BL}}^{2}\bar{\mathring{\Phi}}_{\mathrm{ORG}}\right]=\sum_{\ell,m,n}\frac{\bar{C}_{\ell mn}}{32\sqrt{2\pi}}\Delta_{\mathrm{BL}}^{2}\left(\partial_{r}-i\frac{\bar{K}_{mn}}{\Delta_{\mathrm{BL}}}\right)^{4}\left[\Delta_{\mathrm{BL}}^{2}\,{}_{+2}\bar{R}_{\ell mn}(r)\,{}_{+2}S_{\ell mn}(\theta)e^{-im\phi+i\bar{\omega}_{mn}t}\right]$ (48)
@@
$\frac{1}{32}\Delta_{\mathrm{BL}}^{2}\left(\hat{\mathcal{D}}_{0}^{\dagger}\right)^{4}\left[\Delta_{\mathrm{BL}}^{2}\bar{\mathring{\Phi}}_{\mathrm{ORG}}\right]=\sum_{\ell,m,n}(-1)^{m}\frac{\bar{C}_{\ell-m-n}}{32\sqrt{2\pi}}\Delta_{\mathrm{BL}}^{2}\left(\hat{\mathcal{D}}_{mn}^{\dagger}\right)^{4}\left[\Delta_{\mathrm{BL}}^{2}\,{}_{+2}\bar{R}_{\ell mn}(r)\,{}_{-2}S_{\ell mn}(\theta)e^{im\phi-i\omega_{mn}t}\right]$ (49)
When considering Eq. (38), we obtain
${}_{-2}Z_{\ell mn}\,{}_{-2}R_{\ell mn}(r)=\frac{(-1)^{m}}{32}\bar{C}_{\ell-m-n}\Delta_{\mathrm{BL}}^{2}\left(\hat{\mathcal{D}}_{mn}^{\dagger}\right)^{4}\left[\Delta_{\mathrm{BL}}^{2}\,{}_{+2}R_{\ell mn}(r)\right]\,.$ (50)
The asymptotic behavior of ${}_{s}R_{\ell mn}$ is as follows Teukolsky and Press (1974); Cook and Zalutskiy (2014):
$\lim_{r^{*}\rightarrow\infty}{}_{s}R_{\ell mn}(r^{*})\sim\begin{cases}\frac{e^{-i\omega r^{*}}}{r}&:\text{ ingoing wave }\\ \frac{e^{i\omega r^{*}}}{r^{2s+1}}&:\text{ outgoing wave. }\end{cases}$ (51)
$\lim_{r^{*}\rightarrow-\infty}{}_{s}R_{\ell mn}\left(r^{*}\right)\sim\begin{cases}e^{ik_{mn}r^{*}}&:\text{ ingoing (out of }\mathrm{BH})\\ \Delta_{\mathrm{BL}}^{-s}e^{-ik_{mn}r^{*}}&:\text{ outgoing (into BH) }\end{cases}$ (52)
@@
$r^{*}=r+\frac{2M}{r_{+}-r_{-}}\left[r_{+}\ln\left(\frac{r-r_{+}}{2M}\right)-r_{-}\ln\left(\frac{r-r_{-}}{2M}\right)\right]$ (53)
If $f(r)$ is a solution to the radial Teukolsky equation Eq. (41) with spin $s$, then $\Delta_{\mathrm{BL}}^{s}\bar{f}(r)$ is a solution to the radial Teukolsky equation with spin $-s$. For generic modes with $\omega_{mn}\neq 0$, the asymptotic behavior suggests that this relation swaps physical (retarded) boundary conditions with unphysical (advanced) ones. We denoted the relationship as
${}_{-s}R^{\#}_{\ell mn}\equiv\Delta_{\mathrm{BL}}^{s}\,{}_{s}\bar{R}_{\ell mn}.$ (54)
@@
${}_{-2}Z_{\ell mn}\,{}_{-2}R_{\ell mn}(r)=\frac{(-1)^{m}}{32}\bar{C}_{\ell-m-n}\Delta_{\mathrm{BL}}^{2}\left(\hat{\mathcal{D}}_{mn}^{\dagger}\right)^{4}\left[{}_{-2}R^{\#}_{\ell mn}\right]\,.$ (55)
when we insert the asymptotic behavior into Eq. (55), we can find
$C_{\ell mn}=\frac{2}{\omega_{mn}^{4}}(-1)^{\ell+m}{}_{-2}Z_{\ell mn}$ (56)

### v3

@@ 2 Metric Reconstruction
    ^$\begin{split}\hat{\mathcal{E}}_{ab}(h)=\frac{1}{2}\left[-\nabla^{c}\nabla_{c}h_{ab}-\nabla_{a}\nabla_{b}h_{c}{}^{c}+2\nabla^{c}\nabla_{(a}h_{b)c}+g_{ab}\left(\nabla^{c}\nabla_{c}h_{d}{}^{d}-\nabla^{c}\nabla^{d}h_{cd}\right)\right]\,.\end{split}$ (A.17)
@@
  • The Teukolsky operator $\hat{\mathcal{O}}$, which can be written as [16]
    $\hat{\mathcal{O}}_{0}=$ $\left(\hat{D}-3\epsilon+\bar{\epsilon}-4\rho-\bar{\rho}\right)\left(\hat{\Delta}-4\gamma+\mu\right)-\left(\hat{\delta}+\bar{\varpi}-\bar{\alpha}-3\beta-4\tau\right)\left(\hat{\bar{\delta}}+\varpi-4\alpha\right)-3\Psi_{2}\,,$ (A.18)
@@
    $\hat{\mathcal{O}}_{0}\Psi_{0}^{(1)}=8\pi T_{0}\,,\,\hat{\mathcal{O}}_{4}\Psi_{4}^{(1)}=8\pi T_{4}\,.$ (A.19)
    Their separable form, the Teukolsky master equation, can be written as
    $\hat{\mathcal{O}}_{s}\psi_{s}=8\pi\Sigma T_{s}\,,$ (A.20)
@@
    $\psi_{-2}=\rho^{-4}\Psi_{4}^{(1)}\,,\hat{\mathcal{O}}_{-2}=2\Sigma\rho^{-4}\hat{\mathcal{O}}_{4}\rho^{4}\,,T_{-2}=2\rho^{-4}T_{4}\,;$ (A.21)
@@
    $\psi_{+2}=\Psi_{0}^{(1)},\,\hat{\mathcal{O}}_{+2}=2\Sigma\hat{\mathcal{O}}_{0}\,,T_{+2}=2T_{0}\,.$ (A.22)
  • The operator $\hat{\mathcal{S}}$, which acts on an energy-momentum tensor $T_{ab}$ to yield the source $T_{4}$ of the Teukolsky equation, (Eq. (2.15) in Ref. [16])
    $\hat{\mathcal{S}}^{ab}=$ $\left(\hat{\Delta}+3\gamma-\bar{\gamma}+4\mu+\bar{\mu}\right)\left[(\hat{\overline{\delta}}-2\bar{\tau}+2\alpha)n^{(a}\bar{m}^{b)}-(\hat{\Delta}+2\gamma-2\bar{\gamma}+\bar{\mu})\bar{m}^{a}\bar{m}^{b}\right]$ (A.23)
@@
  • The operator $\hat{\mathcal{T}}$, which acts on a metric perturbation $h_{ab}$ to yield the perturbed Weyl scalar $\Psi_{4}^{(1)}$.
    $\hat{\mathcal{T}}^{ab}=$ $-\frac{1}{2}\left\{\left(\hat{\overline{\delta}}-\bar{\tau}+3\alpha+\bar{\beta}\right)\left(\hat{\overline{\delta}}-\bar{\tau}+2\alpha+2\bar{\beta}\right)n^{a}n^{b}\right.$ (A.24)
@@
Figure S1: Relation of operators for $s=-2$.
As pointed out by Wald [93], one can find the following operator identity
$\hat{\mathcal{O}}_{4}\hat{\mathcal{T}}=\hat{\mathcal{S}}\hat{\mathcal{E}}\,,$ (A.25)
@@
^$\hat{\mathcal{E}}_{ab}(h)=2\operatorname{Re}(\hat{\mathcal{T}}^{\dagger}_{ab}\hat{\mathcal{O}}^{\dagger}_{4}\Phi)=0$ (A.26)
@@
$h_{ab}=2\operatorname{Re}\left(\hat{\mathcal{S}}_{ab}^{\dagger}\Phi\right)\,.$ (A.27)
Chrzanowski introduced two radiation gauges: the ingoing radiation gauge (IRG) and the outgoing radiation gauge (ORG) [26]. In this paper, we choose ORG, which can be written as
$h_{ab}n^{b}=0=g^{ab}h_{ab}\,.$ (A.28)
@@
$\mathcal{O}^{\dagger}_{4}=\left(\hat{D}-3\epsilon+\bar{\epsilon}-\bar{\rho}\right)\left(\hat{\Delta}-4\gamma-3\mu\right)-\left(\hat{\delta}-3\beta-\bar{\alpha}+\bar{\varpi}\right)\left(\hat{\bar{\delta}}-4\alpha-3\varpi\right)-3\Psi_{2}\,.$ (A.30)
@@
$\hat{\mathcal{O}}_{4}^{\dagger}\Phi_{\mathrm{ORG}}=\rho^{-4}\mathcal{O}_{0}\left(\rho^{4}\Phi_{\mathrm{ORG}}\right)=0\,,$ (A.31)
and the metric component can be written as [81]
$h_{ab}=$$\left\{n_{(a}\bar{m}_{b)}\left[\left(\hat{\bar{\delta}}+\bar{\beta}-3\alpha+\bar{\tau}+\varpi\right)\left(\hat{\Delta}-4\gamma-3\mu\right)\right.\right.$ (A.32)
@@
where “c.c.” stands for the complex conjugate part of the whole object. If we define $\mathring{\Phi}_{\mathrm{ORG}}=\rho^{4}\Phi_{\mathrm{ORG}}$, then the non-zero metric component can be written as [85, 96]
$h_{ll}\equiv h_{\mu\nu}l^{\mu}l^{\nu}=\hat{\mathcal{H}}_{ll}^{\dagger}\mathring{\Phi}_{\mathrm{ORG}}+\text{c.c.}\,,\,h_{lm}\equiv h_{\mu\nu}l^{\mu}m^{\nu}=\hat{\mathcal{H}}_{lm}^{\dagger}\mathring{\Phi}_{\mathrm{ORG}},$ (A.33)
@@
^$\hat{\mathcal{H}}_{ll}^{\dagger}$ $=-\rho^{-4}\left(\hat{\bar{\delta}}-3\alpha-\bar{\beta}+5\varpi\right)\left(\hat{\bar{\delta}}-4\alpha+\varpi\right)\,,$ (A.34)
@@
While $\mathring{\Phi}_{\mathrm{ORG}}$ does indeed satisfy the vacuum Teukolsky equation with $s=+2$, the field does not correspond to either the Weyl scalar $\Psi^{\mathrm{(1)}}_{0}$ or $\Psi^{\mathrm{(1)}}_{4}$ originating from the reconstructed metric $h_{ab}$. From Fig. S1, it is clear that $\Psi^{\mathrm{(1)}}_{4}$ can be obtained via a fourth-order ordinary differential equation (ODE) along ingoing null rays [87, 96, 89]
$\Psi_{4}^{(1)}=\hat{\mathcal{T}}^{ab}h_{ab}\,.$ (A.35)
We can reduce this circularity condition to what is known as a radial inversion relation [89, 84, 76, 80, 88]
$\Psi_{4}^{(1)}=\frac{1}{2}\left(\hat{\Delta}+3\gamma-\bar{\gamma}\right)\left(\hat{\Delta}+2\gamma-2\bar{\gamma}\right)\left(\hat{\Delta}+\gamma-3\bar{\gamma}\right)\left(\hat{\Delta}-4\bar{\gamma}\right)\bar{\Phi}_{\mathrm{ORG}}\,,$ (A.36)
which can be solved in vacuum algebraically by applying the separation of variables in the next section.
2.1 QNM to Hertz potential
In Boyer-Lindquist (BL) coordinates, the Teukolsky operator $\hat{\mathcal{O}}_{s}$ with spin $s$ in Eq. (A.20) can be written as [16]
@@ 2.1 QNM to Hertz potential
$\displaystyle-2s\left[\frac{M\left(r^{2}-a^{2}\right)}{\Delta_{\mathrm{BL}}}-r-ia\cos\theta\right]\frac{\partial\psi_{s}}{\partial t}+\left(s^{2}\cot^{2}\theta-s\right)\psi_{s}$ (A.37)
@@
$\psi_{s}(t,r,\theta,\phi)=\frac{1}{\sqrt{2\pi}}\sum_{l,m,n}{}_{s}Z_{\ell mn}\,{}_{s}S_{\ell mn}(\theta)\,{}_{s}R_{\ell mn}(r)e^{-i\omega_{mn}t+im\phi}\,,$ (A.38)
@@
$\partial_{z}\left[\left(1-z^{2}\right)\partial_{z}\left[{}_{s}S_{\ell mn}(\theta)\right]\right]+\left[(cz)^{2}-2csz+s+A^{\ell mn}_{s}(c)-\frac{(m+sz)^{2}}{1-z^{2}}\right]{}_{s}S_{\ell mn}(\theta)=0$ (A.39)
where $c=a\omega$, and ${}_{s}S_{\ell mn}(\theta)e^{im\phi}$ represents the spin-weighted spheroidal harmonics, with their phase is consistent with Ref. [23]. To normalize the spin-weighted spheroidal harmonics, we apply the following normalization condition
@@
In the following, we will revisit the procedure for obtaining the Hertz potential corresponding to specific Weyl scalar $\Psi^{\mathrm{(1)}}_{0}$ or $\Psi^{\mathrm{(1)}}_{4}$, which is discussed in Ref. [87, 96]. According to the previously mentioned condition, $\hat{\mathcal{O}}_{+2}\mathring{\Phi}_{\mathrm{ORG}}=0$, we can expand $\mathring{\Phi}_{\mathrm{ORG}}$ in modes
$\mathring{\Phi}_{\mathrm{ORG}}=\frac{1}{\sqrt{2\pi}}\sum_{\ell,m,n}C_{\ell mn}\,{}_{+2}R_{\ell mn}(r)\,{}_{+2}S_{\ell mn}(\theta)e^{im\phi-i\omega_{mn}t}\,,$ (A.43)
where $C_{\ell mn}$ are constant coefficients. In addition, the Hertz potential $\mathring{\Phi}_{\mathrm{ORG}}$ also satisfies Eq. (A.36), which can be simplified using the following relationship
$\left(\hat{\Delta}-m\bar{\gamma}+n\gamma\right)f=\Delta_{\mathrm{BL}}^{(n-m)/2}\bar{\rho}^{-m}\rho^{n}\hat{\Delta}\left[\Delta_{\mathrm{BL}}^{(m-n)/2}\bar{\rho}^{m}\rho^{-n}f\right]\,.$ (A.44)
In this equation, $f=f(t,r,\theta,\phi)$ can represent any function. Then Eq. (A.36) can be written as [85, 96, 87, 88]
$\frac{1}{32}\Delta_{\mathrm{BL}}^{2}\left(\hat{\mathcal{D}}_{0}^{\dagger}\right)^{4}\left[\Delta_{\mathrm{BL}}^{2}\bar{\mathring{\Phi}}_{\mathrm{ORG}}\right]=\rho^{-4}\Psi_{4}^{\mathrm{(1)}}\,.$ (A.45)
@@
$\hat{\mathcal{D}}_{0}^{\dagger}=\partial_{r}-\frac{\left(r^{2}+a^{2}\right)\partial_{t}+a\partial_{\phi}}{\Delta_{\mathrm{BL}}}\,.$ (A.46)
Note that Eq. (A.45) can be solved easily on the mode basis given by Eq. (A.43), using the operator
$\hat{\mathcal{D}}_{mn}^{\dagger}=\partial_{r}+i\frac{K_{mn}}{\Delta_{\mathrm{BL}}}=\partial_{r}+i\frac{\omega_{mn}\left(r^{2}+a^{2}\right)-ma}{\Delta_{\mathrm{BL}}}\,,$ (A.47)
which acts on radial modes. The coefficient can be fixed by considering the asymptotic limit approaching infinity and the horizon. This analysis was previously carried out in Ref. [87, 96] for real $\omega$. Due to our choice of gauge (ORG), the results are similar to those in Section III.C of Ref. [96]. Here, we briefly review and extend the procedure to complex $\omega$.
$\frac{1}{32}\Delta_{\mathrm{BL}}^{2}\left(\hat{\mathcal{D}}_{0}^{\dagger}\right)^{4}\left[\Delta_{\mathrm{BL}}^{2}\bar{\mathring{\Phi}}_{\mathrm{ORG}}\right]=\sum_{\ell,m,n}\frac{\bar{C}_{\ell mn}}{32\sqrt{2\pi}}\Delta_{\mathrm{BL}}^{2}\left(\partial_{r}-i\frac{\bar{K}_{mn}}{\Delta_{\mathrm{BL}}}\right)^{4}\left[\Delta_{\mathrm{BL}}^{2}\,{}_{+2}\bar{R}_{\ell mn}(r)\,{}_{+2}S_{\ell mn}(\theta)e^{-im\phi+i\bar{\omega}_{mn}t}\right]$ (A.48)
@@
$\frac{1}{32}\Delta_{\mathrm{BL}}^{2}\left(\hat{\mathcal{D}}_{0}^{\dagger}\right)^{4}\left[\Delta_{\mathrm{BL}}^{2}\bar{\mathring{\Phi}}_{\mathrm{ORG}}\right]=\sum_{\ell,m,n}(-1)^{m}\frac{\bar{C}_{\ell-m-n}}{32\sqrt{2\pi}}\Delta_{\mathrm{BL}}^{2}\left(\hat{\mathcal{D}}_{mn}^{\dagger}\right)^{4}\left[\Delta_{\mathrm{BL}}^{2}\,{}_{+2}\bar{R}_{\ell mn}(r)\,{}_{-2}S_{\ell mn}(\theta)e^{im\phi-i\omega_{mn}t}\right]$ (A.49)
When considering Eq. (A.38), we obtain
${}_{-2}Z_{\ell mn}\,{}_{-2}R_{\ell mn}(r)=\frac{(-1)^{m}}{32}\bar{C}_{\ell-m-n}\Delta_{\mathrm{BL}}^{2}\left(\hat{\mathcal{D}}_{mn}^{\dagger}\right)^{4}\left[\Delta_{\mathrm{BL}}^{2}\,{}_{+2}R_{\ell mn}(r)\right]\,.$ (A.50)
The asymptotic behavior of ${}_{s}R_{\ell mn}$ is as follows [91, 23]:
$\lim_{r^{*}\rightarrow\infty}{}_{s}R_{\ell mn}(r^{*})\sim\begin{cases}\frac{e^{-i\omega r^{*}}}{r}&:\text{ ingoing wave }\\ \frac{e^{i\omega r^{*}}}{r^{2s+1}}&:\text{ outgoing wave. }\end{cases}$ (A.51)
$\lim_{r^{*}\rightarrow-\infty}{}_{s}R_{\ell mn}\left(r^{*}\right)\sim\begin{cases}e^{ik_{mn}r^{*}}&:\text{ ingoing (out of }\mathrm{BH})\\ \Delta_{\mathrm{BL}}^{-s}e^{-ik_{mn}r^{*}}&:\text{ outgoing (into BH) }\end{cases}$ (A.52)
@@
$r^{*}=r+\frac{2M}{r_{+}-r_{-}}\left[r_{+}\ln\left(\frac{r-r_{+}}{2M}\right)-r_{-}\ln\left(\frac{r-r_{-}}{2M}\right)\right]$ (A.53)
If $f(r)$ is a solution to the radial Teukolsky equation Eq. (A.41) with spin $s$, then $\Delta_{\mathrm{BL}}^{s}\bar{f}(r)$ is a solution to the radial Teukolsky equation with spin $-s$. For generic modes with $\omega_{mn}\neq 0$, the asymptotic behavior suggests that this relation swaps physical (retarded) boundary conditions with unphysical (advanced) ones. We denoted the relationship as
${}_{-s}R^{\#}_{\ell mn}\equiv\Delta_{\mathrm{BL}}^{s}\,{}_{s}\bar{R}_{\ell mn}.$ (A.54)
@@
${}_{-2}Z_{\ell mn}\,{}_{-2}R_{\ell mn}(r)=\frac{(-1)^{m}}{32}\bar{C}_{\ell-m-n}\Delta_{\mathrm{BL}}^{2}\left(\hat{\mathcal{D}}_{mn}^{\dagger}\right)^{4}\left[{}_{-2}R^{\#}_{\ell mn}\right]\,.$ (A.55)
Inserting the asymptotic behavior into Eq. (A.55), we find
$C_{\ell mn}=\frac{2}{\omega_{mn}^{4}}(-1)^{\ell+m}{}_{-2}Z_{\ell mn}$ (A.56)